# Enabling Systematic Generalization in Abstract Spatial Reasoning through Meta-Learning for Compositionality

### v1

@@ 5.1 Consistency across data splits
To ensure that the strong performance of MLC, as reported in Table 1, is not the result of a favorable data split, we train and evaluate the model on three additional, independently generated data splits for each task configuration—resulting in four distinct models per task setup. Detailed descriptions of these data splits are provided in Table 5 in the Appendix. Table 2 summarizes the average accuracy and corresponding standard deviation across all four splits. For the standard three-shot learning task, MLC consistently achieves high accuracy, with a mean of 98.78% and a standard deviation of 1.99%. Similarly, in the systematicity task, the model demonstrates robust generalization, achieving an even higher average accuracy than on the initial data split, with a mean of 86.73% 6.03%.
| Model | Accuracy [%] | Color Acc. [%] | Shape Acc. [%] |
| MLC (3-Shot) | |||
| MLC (Systematicity) | 86.73 6.03 | 99.36 0.70 | 87.55 5.45 |
| - no copy task | |||
| - no primitives | 99.56 0.50 | ||
| - no level-1 compositions |
@@
In conclusion, our experiments highlight the potential of MLC beyond linguistic tasks. The presented results demonstrate that MLC enables systematic generalization to novel transformation compositions, significantly outperforming current state-of-the-art LLMs.
@@ 6 Related work
Meta-learning aims to improve a model’s ability to adapt to novel tasks by leveraging experience over multiple training episodes (Thrun & Pratt, 1998; Hospedales et al., 2022). It has been successfully applied to various tasks, such as few-shot learning (Mishra et al., 2018), continual learning (Javed & White, 2019; Lee et al., 2023; Irie et al., 2025), and reinforcement learning (Duan et al., 2016; Wang et al., 2017; Mishra et al., 2018). Related to our work, meta-learning has been used to improve systematic generalization. Lake & Baroni (2018) showed that traditional sequence-to-sequence models struggle with compositional skills, but incorporating meta-learning can significantly improve performance (Lake, 2019; Conklin et al., 2021). Recent work argues that giving models the opportunity to practice skills via meta-learning is crucial for addressing challenges such as systematic generalization, among others (Irie & Lake, 2024). Our method builds on meta-learning strategies inspired by Lake & Baroni (2023), extending them to the domain of abstract spatial reasoning.
@@
The abstraction and reasoning corpus (ARC) (Chollet, 2019) is a benchmark designed to evaluate a model’s capacity to generalize to novel scenarios with limited to no prior knowledge. Based on a set of few-shot examples, models are required to infer transformations of abstract objects or patterns within two-dimensional grids. Unlike ARC, which encompasses a broad range of complex transformations, our work deliberately narrows the scope to the five fundamental geometric transformations described in Section 3.1, focusing instead on the aspect of systematicity. Several ARC variants have been proposed, including 1D-ARC (Xu et al., 2023), Mini-ARC (Kim et al., 2022), ConceptARC (Moskvichev et al., 2023) and MC-LARC (Shin et al., 2024). However, our dataset, SYGAR, is the first to focus on the aspect of systematicity in abstract spatial reasoning.
@@ 7 Conclusion
In this work, we extend the meta-learning for compositionality framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. To this end, we introduce SYGAR—a novel dataset designed to evaluate systematicity in this field. Our experiments demonstrate that models trained via MLC can systematically generalize to novel compositions of geometric transformations. Moreover, MLC consistently outperforms current state-of-the-art LLMs, which do not exhibit comparable systematic behavior. Our findings suggest that MLC presents a promising direction for enabling systematic generalization in language models across diverse domains.
@@
To ensure the reproducibility of our work, we make all code publicly available at: https://github.com/mainlp/SYGAR. This enables users to reproduce the data described in Section 3.1 and train models via MLC for the task, as outlined in Section 3.2. Details about the training procedures and hyperparameters are provided in Section 3.2 and Appendix B. Additionally, we include an exemplary subset of input queries and corresponding API responses from the evaluated LLMs as part of the supplementary material. Specifics on prompts, model versions, and decoding parameters are given in Appendix C.2. Further details about the datasets can be found in Section 3.1, Section 4.1, and Appendix A. Finally, Appendix B.2 outlines the software and computational resources used for model training.
@@ Appendix A Dataset
In this work, we present SYGAR, a dataset designed to study systematicity in abstract spatial reasoning. As outlined in Section 3.1, SYGAR evaluates a model’s capacity to systematically generalize learned geometric transformations (e.g., translation, rotation) of two-dimensional objects to novel compositions of these transformations (e.g., translation+rotation). The subsequent sections offer a detailed description of the dataset, including formal definitions of the grid-based environment and the set of transformations it includes.
@@ A.1 Grid setup
[Grid & Object] Let be a matrix, referred to as a grid, where each element . The value represents a background cell, and values represent object colors.
@@ A.2 Geometric transformations
[Rotation] Let be an object in a grid , and let be the top-left coordinate of the object’s bounding box. Let be the rotation angle.
For each , let the relative offset be:
Let the rotated offset be given by:
Then the rotated object is:
@@ A.3 Dataset generation
To generate episodes that comprise primitive transformations, level-1 transformation compositions, and level-2 transformation compositions, we developed a script that systematically generates the corresponding input-output grid pairs for each transformation. The complete code repository for data generation is publicly available at: https://github.com/mainlp/SYGAR. In the following, we provide a brief overview of the procedure used to generate input-output grid pairs for each sample within an episode. As detailed in Section 3.1 and Appendix A.2, we consider five basic geometric transformations, along with three types of transformation indicators: shape-based, color-based, and neighbor-based. These allow us to define a total of ten distinct transformation triplets, each mapping the indicators to corresponding transformations (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). For each episode, a transformation triplet is uniformly sampled from this set to define the visual interpretation grammar of the episode. Once the transformations are determined, we randomly assign a specific shape for the shape-based transformation, a specific color for the color-based transformation, and an indicator object for the neighbor-based transformation. Importantly, the indicator object is constrained to neither share the shape associated with the shape-based transformation nor the color linked to the color-based transformation.
@@ Appendix B Training details
| Parameter | Value |
| number layers in decoder | 3 |
| number layers in decoder | 3 |
| number of attention heads | 8 |
| hidden dimension | 128 |
| feedforward hidden size | 768 |
| learning rate | 0.01 |
| learning rate after training | |
| dropout | 0.0 |
| weight decay | 0.01 |
| noise probability | 0.001 |
| gradient accumulation over batches | 2 |
| background patch loss weight | 0.2 |
@@ C.2 Model information
As described in Section 4.2, we evaluate three different LLMs in addition to our model trained via MLC. Specifically, we assess the performance of GPT-4o (Achiam et al., 2023) (version gpt-4o-2024-08-06111https://platform.openai.com/docs/models/gpt-4o), o3-mini (OpenAI, 2025) (version o3-mini-2025-01-31222https://platform.openai.com/docs/models/o3-mini), and Gemini 2.0 Flash (DeepMind, 2024) (version gemini-2.0-flash-001333https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash). All models are accessed via their respective batch APIs, enabling us to process multiple samples per request. Unless otherwise specified, we employ the default API settings. For GPT-4o and o3-mini, this corresponds to a temperature and top_p value of .444https://platform.openai.com/docs/api-reference/chat/create Due to financial constraints, the o3-mini model is configured with a “low” reasoning effort. For Gemini 2.0 Flash, the provider does not disclose default parameter settings.
@@ Appendix D Additional results
In this section, we present additional results for the experiments conducted in this study. First, we present additional qualitative results related to the model predictions on the standard few-shot learning and the systematicity task. Figures 4 through 6 illustrate representative episodes from the standard few-shot learning task. Model predictions are shown adjacent to each query, with results for GPT-4o and Gemini 2.0 Flash corresponding to text-only prompts. Across all three episodes, the model trained using MLC consistently predicts the correct output grid. In contrast, GPT-4o and Gemini 2.0 Flash frequently fail to identify the correct transformation—either misrepresenting the shape of the transformed object or incorrectly predicting its final position. Notably, o3-mini successfully predicts the correct output for the episodes in Figures 5 and 6, but fails on the example in Figure 4. Figures 7 and 8 highlight episodes from the systematicity task. As shown, all baseline models fail to produce accurate transformations, often misplacing the transformed object within the grid. In contrast, the model trained via MLC consistently predicts the correct transformation.
| Model | Valid Responses (3-Shot) | Valid Responses (Systematicity) |
| GPT-4o (text-only) | ||
| GPT-4o (text+image) | ||
| Gemini 2.0 Flash (text-only) | ||
| Gemini 2.0 Flash (text+image) | ||
| o3-mini | ||
| MLC |
@@
As outlined in Section C.2, LLMs we evaluate are instructed to present their final output grid predictions using the keyword “output:”, followed by a two-dimensional array of size in the format “[[]]”. Responses that do not conform to this expected pattern are excluded from subsequent analyses and are not included in accuracy calculations. Table 4 provides an overview of the proportion of valid responses for each model. In the standard few-shot learning setting, all models demonstrate very high valid response rates, exceeding 99%. However, in the systematicity task, a slight decrease in valid responses is observed for Gemini 2.0 Flash when additional visual input (text+image) is introduced, with the rate falling to 94.09%. More significantly, GPT-4o exhibits a notable drop in valid response rate to 77.24% under multimodal conditions. We hypothesize that this decline may be attributed to the increased context length resulting from the additional image input.
@@
Training on static data.
In addition to the model trained via MLC on a stream of dynamically changing visual interpretation grammars, as described in Section 3.2, we adopt the approach of Lake (2019) and train a transformer-based encoder-decoder on a dataset governed by a fixed visual grammar (referred to as basic seq2seq). Instead of episodes containing multiple study examples and queries, this dataset comprises individual input-output grid pairs, where the objective is to predict the output grid corresponding to a given input grid. This more closely resembles a standard training approach.
We construct a dataset of 1,300 grid pairs, partitioned into 1,260 training samples, 20 validation samples, and 20 test samples. All examples conform to a fixed visual interpretation grammar, meaning that a set of fixed visual indicators consistently maps to specific transformations. Samples represent primitive transformations, as well as level-1 and level-2 transformation compositions. As with our other experiments, the test set includes level-2 transformation compositions that were not observed during training—only their constituent components and level-1 compositions were seen during training. For instance, the test set might include transformations composed of shape-based downward translation, color-based horizontal reflection, and neighbor-based upward extension. However, only their decomposed elements have been shown during training.
The model is trained for 200 episodes using the parameters specified in Section B. While it successfully fits the training data (accuracy of over 99%), it fails to generalize to the test set, achieving a test accuracy of 0.0%.

### v2

@@ 5.1 Consistency across data splits
To ensure that the strong performance of MLC, as reported in Table 1, is not the result of a favorable data split, we train and evaluate the model on three additional, independently generated data splits for each task configuration—resulting in four distinct models per task setup. Detailed descriptions of these data splits are provided in Table 5 in the Appendix. Table 2 summarizes the average accuracy and corresponding standard deviation across all four splits. For the standard three-shot learning task, MLC consistently achieves high accuracy, with a mean of 98.78% and a standard deviation of 1.99%. Similarly, for the systematicity task, the model demonstrates robust generalization, achieving an even higher average accuracy than on the initial data split, with a mean of 86.73%.
@@
In conclusion, our experiments highlight the potential of MLC beyond linguistic tasks, demonstrating its capacity for systematic generalization in abstract spatial reasoning.
@@ 6 Related work
Meta-learning aims to improve a model’s ability to adapt to novel tasks by leveraging experience over multiple training episodes (Thrun & Pratt, 1998; Hospedales et al., 2022). It has been successfully applied to various tasks, such as few-shot learning (Mishra et al., 2018), continual learning (Javed & White, 2019; Lee et al., 2023; Irie et al., 2025), and reinforcement learning (Duan et al., 2016; Wang et al., 2017; Mishra et al., 2018). Related to our work, meta-learning has been used to improve systematic generalization. Lake & Baroni (2018) showed that traditional sequence-to-sequence models struggle with compositional skills, but incorporating meta-learning can significantly improve performance (Lake, 2019; Conklin et al., 2021). Recent work argues that giving models the opportunity to practice skills via meta-learning is crucial for addressing challenges such as systematic generalization, among others (Irie et al., 2025). Our method builds on meta-learning strategies inspired by Lake & Baroni (2023), extending them to the domain of abstract spatial reasoning.
@@
The abstraction and reasoning corpus (ARC) (Chollet, 2019) is a benchmark designed to evaluate a model’s capacity to generalize to novel scenarios with limited to no prior knowledge. Based on a set of few-shot examples, models are required to infer transformations of abstract objects or patterns within two-dimensional grids. Unlike ARC, which encompasses a broad range of complex transformations, our work deliberately narrows the scope to the five fundamental geometric transformations described in Section 3.1, focusing instead on the aspect of systematicity. Several ARC variants have been proposed, including 1D-ARC (Xu et al., 2024), Mini-ARC (Kim et al., 2022), ConceptARC (Moskvichev et al., 2023) and MC-LARC (Shin et al., 2024). However, to the best of our knowledge, Compositional-ARC is the first to focus on compositional generalization.
@@ 7 Conclusion
In this work, we extend the meta-learning for compositionality framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. To this end, we introduce Compositional-ARC—a novel dataset designed to evaluate systematicity in this field. Our experiments demonstrate that models trained via MLC can systematically generalize to novel compositions of geometric transformations. Moreover, a small MLC model outperforms state-of-the-art general-purpose LLMs on Compositional-ARC, and performs on par with domain-specific LLMs trained via test-time training. Our findings suggest that MLC presents a promising direction for enabling systematic generalization in language models across diverse domains.
@@
To ensure the reproducibility of our work, we make all code publicly available at: https://github.com/mainlp/C-ARC. This enables users to reproduce the data described in Section 3.1 and train models via MLC for the task, as outlined in Section 3.2. Details about the training procedures and hyperparameters are provided in Section 3.2 and Appendix B. Specifics on prompts, model versions, and decoding parameters are given in Appendix C.2. Further details about the datasets can be found in Section 3.1, Section 4.1, and Appendix A. Finally, Appendix B.2 outlines the software and computational resources used for model training.
@@ Appendix A Dataset
In this work, we present Compositional-ARC, a dataset designed to study systematicity in abstract spatial reasoning. As outlined in Section 3.1, Compositional-ARC evaluates a model’s capacity to systematically generalize learned geometric transformations (e.g., translation, rotation) of two-dimensional objects to novel compositions of these transformations (e.g., translation+rotation). The subsequent sections offer a detailed description of the dataset, including formal definitions of the grid-based environment and the set of transformations it includes.
@@ A.1 Grid setup
[Grid & Object] Let be a matrix with rows and columns , referred to as a grid, where each element . The value represents a background cell, and values represent object colors.
@@ A.2 Geometric transformations
[Rotation] Let be a set of grid cells with row–column coordinates . Let and . We set the pivot as the top-left of the bounding box.
For each , we specify the offset from the pivot as:
We define a rotation by as:
where is clockwise and is counterclockwise under the row-down convention.
Given a rotation, either clockwise or counterclockwise, the rotated object is:
@@ A.3 Dataset generation
To generate episodes that comprise primitive transformations, level-1 transformation compositions, and level-2 transformation compositions, we developed a script that systematically generates the corresponding input-output grid pairs for each transformation. The complete code repository for data generation is publicly available at: https://github.com/mainlp/C-ARC. In the following, we provide a brief overview of the procedure used to generate input-output grid pairs for each sample within an episode. As detailed in Section 3.1 and Appendix A.2, we consider five basic geometric transformations, along with three types of transformation indicators: shape-based, color-based, and neighbor-based. These allow us to define a total of ten distinct transformation triplets, each mapping the indicators to corresponding transformations (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). For each episode, a transformation triplet is uniformly sampled from this set to define the visual interpretation grammar of the episode. Once the transformations are determined, we randomly assign a specific shape for the shape-based transformation, a specific color for the color-based transformation, and an indicator object for the neighbor-based transformation. Importantly, the indicator object is constrained to neither share the shape associated with the shape-based transformation nor the color linked to the color-based transformation.
@@ Appendix B Training details
| Parameter | Value | Parameter | Value | |
| number layers in decoder | 3 | learning rate after training | ||
| number layers in decoder | 3 | dropout | 0.0 | |
| number of attention heads | 8 | weight decay | 0.01 | |
| hidden dimension | 128 | noise probability | 0.001 | |
| feedforward hidden size | 768 | gradient accumulation over batches | 2 | |
| learning rate | 0.01 | background patch loss weight | 0.2 |
@@ C.2 Model information
General-purpose LLMs.
As described in Section 4.2, we evaluate three different general-purpose LLMs on Compositional-ARC. Specifically, we assess the performance of o3-mini (OpenAI, 2025) (version o3-mini-2025-01-31222https://platform.openai.com/docs/models/o3-mini), GPT-4o (Achiam et al., 2023) (version gpt-4o-2024-08-06333https://platform.openai.com/docs/models/gpt-4o), and Gemini 2.0 Flash (DeepMind, 2024) (version gemini-2.0-flash-001444https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash). All models are accessed via their respective batch APIs, enabling us to process multiple samples per request. Unless otherwise specified, we employ the default API settings. For GPT-4o and o3-mini, this corresponds to a temperature and top_p value of .555https://platform.openai.com/docs/api-reference/chat/create Due to financial constraints, the o3-mini model is configured with a “low” reasoning effort. For Gemini 2.0 Flash, the provider does not disclose default parameter settings.
| Model | Valid Responses (3-Shot) | Valid Responses (Systematicity) |
| GPT-4o | 99.95% | 99.40% |
| + image | 99.80% | 77.24% |
| Gemini 2.0 Flash | 99.92% | 99.74% |
| + image | 99.51% | 94.09% |
| o3-mini (low) | 100% | 100% |
| Llama-3.2-3B-ReARC | 100% | 100% |
| + test-time training | - | 100% |
| Mistral-NeMO-Minitron-8B-Full | 100% | 100% |
| + test-time training | - | 100% |
| MLC (ours) | 100% | 100% |
@@
Domain-specific LLMs.
As mentioned in Section 4.2, we also evaluate two LLMs proposed by Franzen et al. (2024) that are specifically tailored to ARC-style data: (i) Llama-3.2-3B-ReARC (version Llama-3.2-3B-ARChitects-ReArc-bnb-4bit666https://huggingface.co/da-fr/Llama-3.2-3B-ARChitects-ReArc-bnb-4bit) and (ii) Mistral-NeMO-Minitron-8B-Full (version Mistral-NeMo-Minitron-8B-ARChitects-Full-bnb-4bit777https://huggingface.co/da-fr/Mistral-NeMo-Minitron-8B-ARChitects-Full-bnb-4bit). We use the original code888https://github.com/da-fr/arc-prize-2024 provided by the authors to run their models on Compositional-ARC, with default parameters. This means that the models perform augmented inference on the test set with rotations and transpositions over all symmetries, in addition to color permutations and example shuffling. Candidate pruning is further applied with a minimum probability of 0.1. For models evaluated with test-time training, we follow the authors’ one-epoch LoRA adaptation on the study examples of the test data repeated 48 times with the same augmentations described before. LoRA targets the attention and MLP modules, as well as the embeddings, with , , and dropout set to 0. The models are trained with a batch size of 16, gradient accumulation set to 1, a cosine learning rate of (with for embeddings), and a warmup ratio of 0.25. The resulting weights are then used for inference with the same default settings as described earlier.
@@ Appendix D Additional results
In this section, we present additional results for the experiments conducted in this study. First, we present additional qualitative results related to the model predictions on the standard few-shot learning and the systematicity task. Figures 4 through 6 illustrate representative episodes from the standard few-shot learning task. Model predictions are shown adjacent to each query, with results for GPT-4o and Gemini 2.0 Flash corresponding to text-only prompts. Across all three episodes, the model trained using MLC consistently predicts the correct output grid. In contrast, GPT-4o and Gemini 2.0 Flash frequently fail to identify the correct transformation—either misrepresenting the shape of the transformed object or incorrectly predicting its final position. Notably, o3-mini successfully predicts the correct output for the episodes in Figures 5 and 6, but fails on the example in Figure 4. Figures 7 and 8 highlight episodes from the systematicity task. As shown, all general-purpose LLMs fail to produce accurate transformations, often misplacing the transformed object within the grid. In contrast, the model trained via MLC consistently predicts the correct transformation.
@@
As outlined in Section C.2, the general-purpose LLMs we evaluate are instructed to present their final output grid predictions using the keyword “output:”, followed by a two-dimensional array of size in the format “[[]]”. Responses that do not conform to this expected pattern are excluded from subsequent analyses and are not included in accuracy calculations. Table 4 provides an overview of the proportion of valid responses for each model. In the standard few-shot learning setting, all models demonstrate very high valid response rates, exceeding 99%. However, in the systematicity task, a slight decrease in valid responses is observed for Gemini 2.0 Flash when additional visual input (text+image) is introduced, with the rate falling to 94.09%. More significantly, GPT-4o exhibits a notable drop in valid response rate to 77.24% under multimodal conditions. We hypothesize that this decline may be attributed to the increased context length resulting from the additional image input.
Training on static data.
In addition to the model trained via MLC on a stream of dynamically changing visual interpretation grammars, as described in Section 3.2, we adopt the approach of Lake (2019) and train a transformer-based encoder-decoder on a dataset governed by a fixed visual grammar (referred to as basic seq2seq). This means that the indicator-transformation mappings are static across the whole dataset. For instance, if yellow object translates one step downward, then this applies to all data samples across the dataset. Instead of episodes with few-shot examples, this dataset comprises individual input-output grid pairs, where the objective is to predict the output grid corresponding to a given input grid. This more closely resembles a standard training approach.
We construct a dataset of 1,300 grid pairs, partitioned into 1,260 training samples, 20 validation samples, and 20 test samples. Samples represent primitive transformations, as well as level-1 and level-2 transformation compositions. As with our other experiments, the test set includes level-2 transformation compositions that were not observed during training—only their constituent components and level-1 compositions were seen during training. For instance, the test set might include transformations composed of shape-based downward translation, color-based horizontal reflection, and neighbor-based upward extension. However, only their decomposed elements have been shown during training.
The model is trained for 200 epochs on the dataset using the parameters specified in Section B. While it successfully fits the training data (with an accuracy of over 99%), it fails to generalize to the out-of-distribution test set, achieving a test accuracy of 0.0%. This demonstrates that traditional model training, sample by sample, does not encourage systematic generalization to unseen compositions. Instead, systematicity requires a training procedure with examples over dynamically varying interpretation grammars, as described in Section 3.2.
Appendix E Use of AI assistants
We used GitHub Copilot for parts of the project’s code, and ChatGPT for minor language revisions.